OpenStudy (anonymous):

Let t= 5 What is the value of 5t - 5?

OpenStudy (lizz123):

Rewrite the equation with t.

OpenStudy (lizz123):

It would be 5[t=5]-5

OpenStudy (anonymous):

Kk

OpenStudy (lizz123):

What would you do if two numbers are together? Multiply or divide.

OpenStudy (anonymous):

5(5)-5=25-5=? Answer=...

OpenStudy (anonymous):

Divide?

OpenStudy (anonymous):

your answer is 20

OpenStudy (anonymous):

You multiply

OpenStudy (lizz123):

no when two variables are togther you multiply.

OpenStudy (anonymous):

Thanks

OpenStudy (lizz123):

So what is 5 times 5?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

25

OpenStudy (lizz123):

and the subtract that by 5.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

20

OpenStudy (anonymous):

yes

OpenStudy (lizz123):

Therefore, 20 is your answer. Do you need anymore help?

OpenStudy (anonymous):

Nope now since u explained it to me, for my other answers which I was 2 more I know how to do them .

OpenStudy (anonymous):

*their is to more

OpenStudy (lizz123):

sure what are the other questions

OpenStudy (anonymous):

Ok, thanks Let n = 5. What is the value of 10 + 15 divided by n - 2?

OpenStudy (anonymous):

But I have the answer to that its easy

OpenStudy (lizz123):

So rewrite the equation: 10+15/5 [which is n] -2 So you need to know PEMDAS or Please Excuse my Dear Aunt Sally P is parenthesis E is exponents M is multiply D is divide A is add S is subtract So what's first + or / [divide]? You + [add] first. So you add 10+15=? So 10+15 = 25 So now rewrite the new equation which is: 25/5-2 Now, what is first using PEMDAS or Please Excuse My Dear Aunt Sally / [divide] or - [subtract]? Divide is first usig PEMDAS or Please Excuse My Dear Aunt Sally. 25/5=? 25/5=5 Now rewrite the new equation. 5-2 So now your last step to subtract. What is 5-2=? The answer you get is 3. Therefore, 3 is your answer.